OpenStudy (anonymous):

If the zeros of g ( x) are x = −2and x = 3, what are the zeros of y = 2 g( x+1) ?

OpenStudy (anonymous):

-0.5 and 2

OpenStudy (anonymous):

sorry it should be -1.5 and 1

OpenStudy (anonymous):

professor Hatcrews is everywhere :D

OpenStudy (anonymous):

You should show the steps @HatcrewS

OpenStudy (anonymous):

g(x+1) : x= -1 and x= 2 2g(x+1) : x= -1 and x= 2 (scaling of y does not affect x)

OpenStudy (anonymous):

g(x+1) : x= -3 and x= 2 (shift parallel to x-axis by -1 unit) 2g(x+1) : x= -3 and x= 2 (scaling of y does not affect x)

OpenStudy (anonymous):

sorry, the right answer is x=-3 x=2 g(-2)=0 g(3)=0 x+1=-2 x=-3 x+1=3 x=2